how do I convert these repeating decimals, 0.7, 2.37 and -4.34

Answer by Alan3354(69443)

how do I convert these repeating decimals,
0.7
If it's 0.777...
n = 0.777...
10n = 7.777...
-1n =-0.777...
9n = 7
n = 7/9
------------
2.37
n = 0.373737...
100n = 37.373737...
--> n = 37/99 + 2
------------
and -4.34
Same as 2.37...